OpenStudy (anonymous):

which of the following is the equation of a line perpendicular to the line -3/2x + 4, passing through the points (3,9)

OpenStudy (anonymous):

We know it is perpendicular to a line of gradient -3/2, so what's the gradient of the new line?

OpenStudy (anonymous):

Well in my notes it says the the slope of 2 perpendicular lines are negatice reciprocals that's really what I'm confused about , I'm not really sure what gradient means

OpenStudy (anonymous):

*negative

OpenStudy (anonymous):

Exactly, it's a negative reciprocal. Do you know what a reciprocal is?

OpenStudy (anonymous):

Gradient is just how steep the line is

OpenStudy (anonymous):



OpenStudy (anonymous):

Ohh okay and yeah doesn't reciprocal just mean equal ?

OpenStudy (anonymous):

Nope, it means you turn the fraction upside down. As a few examples, what are the reciprocals of these: 1. 3/2 2. 1/2 3. 6/5 4. 8

OpenStudy (anonymous):

Ohh okay so in this problem do I flip the -3/2x to -2/3 ?

OpenStudy (anonymous):

That gets the reciprocal, but what does the definition of perpendicular say?

OpenStudy (anonymous):

I'm sure I sound really stupid but I'm not actually sure

OpenStudy (anonymous):

It just says that two lines with slopes m1 ad m2 are perpendicular if and only if their slopes are negative reciprocals

OpenStudy (anonymous):

Don't worry about how you sound, you're here to learn and there's nothing stupid about that :)

OpenStudy (anonymous):

Yes, the negative reciprocal

OpenStudy (anonymous):

So -3/2 goes to?

OpenStudy (anonymous):

-2/3 ? and thank you

OpenStudy (anonymous):

m=-3/2, what is -1/m?

OpenStudy (anonymous):

Are you asking what the reciprocal of -1/m is ?

OpenStudy (anonymous):

Nope, sorry I should've been more clear. The gradient is often referred to as m and in this case m=-3/2. A line perpendicular to a line of gradient m, has gradient -1/m (the negative reciprocal). So what I'm asking you to do is: \[-\frac{1}{-\frac{3}{2}} \]

OpenStudy (anonymous):

Sorry I am just so confused also I left out a few things so I'll retype the question Which of the following is the equation of a line perpendicular to the line y= -3/2x + 4, passing through the points (3,9) A. -2x - 3y = -21 B. -2x +3y = 21 C. 2x + 3y = 21 D. 2x-3y = 21

OpenStudy (anonymous):

The negatives cancel, and 1/(3/2)=2/3 so that's the gradient of our new line. The equation of a straight line is given as (y-y1)=m(x-x1) where m is the gradient and (x1, y1) is a point on this line. We have m and a point (3,9), so we get: y-9=(2/3)(x-3) 3(y-9)=2(x-3) 3y-27=2x-6 -2x+3y=21 So B is the answer :)

OpenStudy (anonymous):

I'm going to bed but if you have any questions I'll answer in the morning.

OpenStudy (anonymous):

Thank you so much !